From a normal pack of cards, card is drawn at random. The probability of getting a Jack or a king is

Grade:9

1 Answers

Arun
6 years ago
Dear Devansh
 
Probability of jack or king is = 4/52 + 4/52
= 8/52
= 2/13
 
Regards
Arun (askIITians forum expert)